# RadarPLM: Adapting Pretrained Language Models for Marine Radar Target Detection with Preference-aware Loss

### v3

@@ Abstract
Recent advances in pre-trained language models (PLMs) have demonstrated their capabilities in capturing universal knowledge, making them promising applications for radar signal processing. Nevertheless, directly fine-tuning PLMs on radar signals is both computationally expensive and prone to overfitting, particularly in low signal-to-clutter ratio (SCR) environments. In this paper, we propose a novel fine-tuning framework for PLM-based marine radar target detection. First, we design a lightweight adaptation module, enabling parameter-efficient fine-tuning while preserving the pretrained model’s general knowledge. Second, a novel preference-aware loss is developed to selectively optimize different feature patches based on their online evaluated learning values, guiding the model to concentrate on the most generalizable feature patterns during optimization. Extensive experiments on real-world marine radar datasets demonstrate that the proposed finetuning framework achieves an average performance improvement of 9.9% over the standard approach under low SCR conditions. Furthermore, the fine-tuned model, RadarPLM, consistently outperforms state-of-the-art detectors, particularly when training data are limited.
@@ I Introduction
With the growth of global economic integration, maritime transportation has become crucial for international trade and security. Real-time and precise detection of maritime targets is essential for intelligent monitoring systems and various applications such as search and rescue, environmental monitoring, and national security. Radar technology, known for its ability to operate under diverse weather conditions and around the clock, plays a key role in maritime target detection. However, radar systems are challenged by sea clutter, which significantly obscures small targets and complicates detection. The complex nature of sea clutter makes it difficult to distinguish weak targets from sea clutter, posing a persistent challenge in radar target detection. Detecting small maritime targets, including small boats, frogmen, and fairway buoys, remains an active research area, as these targets have small radar cross sections and are difficult to identify in cluttered environments.
The early methods for detecting marine targets are typically developed based on the fundamental concept of constant false alarm rate (CFAR), with specific implementations such as CA-CFAR and GO-CFAR. These methods operate by estimating clutter statistics from reference cells to dynamically adjust detection thresholds. Although computationally efficient, these methods exhibit significant limitations in complex marine environments. Empirical studies reveal that their actual false alarm rates (FAR) often deviate drastically from preset parameters, sometimes exceeding theoretical values by orders of magnitude. This inadequacy stems from their reliance on simplistic statistical assumptions, which fail to hold in heterogeneous sea clutter scenarios characterized by spatially varying distributions and nonstationary behaviors. With the development of modern radars with high range and velocity resolution, approaches leveraging multi-domain features of radar echoes have gained significant attention, utilizing phase, Doppler, and time-frequency domain characteristics to distinguish targets from sea clutter. Many researchers have actively engaged in developing approaches that utilize hand-crafted features and machine learning classifiers [1, 2, 3, 4, 5, 6]. However, these methods rely heavily on domain expertise and hand-crafted heuristics, often struggling to capture high-dimensional and complex patterns in the signal feature.
With the rapid development of deep learning technology, deep learning (DL) and neural networks (NN) have gained huge popularity in the computer vision (CV) society and natural language processing (NLP) society. Moreover, DL and NN approaches are also introduced for small marine target detection. These methods primarily utilize radar signal features to train deep neural networks, eliminating the need for precise mathematical models of clutter and targets. As a result, they somewhat reduce the impact of complex environmental conditions on target detection performance by learning to extract meaningful and distinctive features in a high-dimensional space. Some studies use convolutional neural networks (CNNs) [7, 8, 9, 10, 11, 12, 13], recurrent neural networks (RNNs) [14], and graph neural networks (GNNs) [15, 16, 17]. Among all the methods mentioned above, detectors based on sequence features and RNNs [14] or CNNs [10] offer an excellent trade-off between inference speed and detection performance.
Despite recent progress, compact NN models with small parameter scales and randomly initialized weights often exhibit limited learning capabilities and unstable training processes. These limitations lead to poor detection performance, particularly in scenarios with scarce training data or with significant environmental variability. Extensive recent studies [18, 19, 20, 21, 22, 23, 24] have shown that pre-trained language models (PLMs) can be good starting point for wireless signal and time series modeling, as they are pre-trained on massive sequential datasets. Compared to random initialized weights of small models, this starting point may significantly improve model optimization stability and effectiveness. Moreover, the self-attention mechanism in PLM exhibits structural similarity to Principal Component Analysis (PCA) [18], demonstrating strong generality in extracting key features from high-dimensional sequence data. This insight opens a very promising direction for radar target detection by fine-tuning PLMs, offering the potential for performance improvements.
@@
Although powerful and promising, some challenges still exist. First, the large scale of PLM requires lightweight and efficient fine-tuning strategies, which are essential for practical deployment, especially in resource-constrained settings. Second, for practical radar applications, adapting PLMs to process sequence features in low signal-to-clutter ratio (SCR) conditions often leads to severe overfitting issues. This phenomenon is fundamentally attributed to the intrinsic characteristics of sequence features:
@@
To overcome the above-mentioned issues, including the demand for lightweight fine-tuning and the tendency of overfitting under low SCR conditions, we develop a novel framework named RadarPLM. The proposed framework effectively bridges the gap between PLMs and radar target detection tasks through lightweight adaptation module and selective training. First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens. Next, a radar target detector is constructed based on a PLM, which is fine-tuned through a lightweight adaptation module to efficiently align with the radar detection task. For model optimization, we design a preference-aware loss function by incorporating a lightweight reference model, which is inspired by recent advances [25, 26]. This loss introduces a selective training strategy that dynamically adjusts the learning weights of different feature patches according to their online-evaluated learning values, significantly mitigating model overfitting. Finally, the binary classification head is retrained to further improve detection performance.
Our preliminary study [27] demonstrated that fine-tuning a pre-trained GPT2 model effectively extracts discriminative information from sequence features, achieving state-of-the-art detection performance. However, model overfitting remains a challenge. To address this, we propose the RadarLLM framework, which introduces a lightweight adaptation module and selective training strategy for more reliable and generalizable adaptation.
Here we summarize our key contributions as follows:
@@
    We propose RadarPLM, a novel framework combining a lightweight reference model and a fine-tuned PLM for marine radar target detection. To the best of our knowledge, we are the first to show that the well-known LLMs can be a strong starting point for intelligent radar signal processing.
@@
    For effective fine-tuning, we develop a lightweight adaptation module together with a novel preference-aware loss function, which jointly reduce computational overhead and effectively mitigate the risk of model overfitting.
@@
    Our extensive experiments validate that RadarLLM achieves new state-of-the-art detection performance on popular marine radar datasets under both sufficient and limited training data settings.
@@ II Related Work
II-A Deep Learning Powered Marine Radar Target Detection
@@
II-B Language Model Powered Signal Processing Task
Recently, LLMs such as ChatGPT and DeepSeek have demonstrated exceptional capabilities in natural language understanding, code generation, math problem solving, and human-like text generation [30]. Some studies also demonstrate their potential in the signal processing (SP) task. For example, Zhou et al. [18] propose a unified time series analysis framework by fine-tuning frozen GPT2 [31] and achieving SOTA performance on various datasets. Liu et al. [24] propose a GPT2-empowered channel prediction framework to improve prediction accuracy. Sheng et al. [32] utilize the GPT2 model to develop an effective and robust beam prediction method. Zheng and Dai [21] propose a multi-task PLM framework to satisfy the requirements of different wireless communication tasks. IOT-LLM and Penetrative AI [33, 34] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks. However, recent studies [35] have raised concerns about the universal applicability of PLM in signal processing tasks. PLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks. In marine radar target detection under low SCR conditions, directly fine-tuning a PLM often leads to severe overfitting issues, limiting its parameter transfer capability. To mitigate this limitation, we introduce a preference-aware loss that selectively trains on informative feature patches.
@@ III-B Sequence Feature Extraction and Patching
(a) IP feature
(b) DSE feature
(c) SMS feature
(d) Amp feature
(e) DP feature (Target)
(f) DP feature (Clutter)
@@ III-D1 Methodology
  • High Importance (Unlearned): If $\mathcal{L}_{\theta_{{t}}}$ is much larger than $\alpha\mathcal{L}_{\theta_{{t}}}$, the feature token is identified as containing valuable, yet unlearned patterns.
@@ III-D2 Theoretical Insight of the Effectiveness of Feature Token Importance
A. Problem Definition: Inspired by [26], we begin by formalizing an optimization problem of selecting useful feature patches for model fine-tuning. At each training step $t$, given a candidate set $B_{t}$, the objective is to select a feature patch $(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}$ whose inclusion in the training set $D_{t}$ maximally reduces the generalization loss on the unseen test set $D_{\mathrm{te}}$. This can be expressed as the following optimization problem:
$\arg\min_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}-\log p\left(y^{\text{te }}\mid F^{\text{te }};D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right).$ (16)
@@
$-\log p\!\bigl(y^{\mathrm{te}}\mid F^{\mathrm{te}};D_{t}\!\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\bigr)$
$\quad=-\log\frac{p\!\bigl(y,y^{\mathrm{te}}\mid F,F^{\mathrm{te}},D_{t}\bigr)}{p\!\bigl(y\mid F;D_{t}\bigr)}.$ (17)
@@
$p\!\bigl(y,y^{\mathrm{te}}\mid F,F^{\mathrm{te}},D_{t}\bigr)\\ =p\!\bigl(y\mid F;F^{\mathrm{te}},y^{\mathrm{te}},D_{t}\bigr)\,p\!\bigl(y^{\mathrm{te}}\mid F^{\mathrm{te}};F,D_{t}\bigr)\\ =p\!\bigl(y\mid F;y^{\mathrm{te}},F^{\mathrm{te}},D_{t}\bigr)\,p\!\bigl(y^{\mathrm{te}}\mid F^{\mathrm{te}};D_{t}\bigr).$ (18)
@@
$\log p\!\left(y^{\text{te}}\mid F^{\text{te}};\,D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right)$ (19)
$=\log\frac{p\left(y\mid F;y^{\text{te }},F^{\text{te }},D_{t}\right)p\left(y^{\text{te }}\mid F^{\text{te }},D_{t}\right)}{p\left(y\mid F;D_{t}\right)}$
$\;\propto\;\;\log p\left(y\mid F;y^{\mathrm{te}},F^{\mathrm{te}},D_{t}\right)-\log p\left(y\mid F;D_{t}\right)$
$\;\propto\;\;\mathcal{L}\!\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\!\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t},D_{\text{te}}\right].$
In summary, for a model trained on $D_{t}$, the process of identifying the feature patch that minimizes the loss on the test samples in Eq.(16) can be approximated by the following objective:
$\arg\max_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{test}}\right].$ (20)
As exact Bayesian inference is intractable in neural networks, we fit the models with SGD instead. Here, $\mathcal{L}[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}]$ denotes the training loss on the patch using the current model trained on $D_{t}$, whereas $\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{test}}\right]$ represents the excepted irreducible holdout loss obtained if the model were trained on both the training set and the test set.
C. Conclusion: Eq.(20) reveals that the gap between the training loss and the irreducible holdout loss quantifies the reducible component of the generalization error of each feature patch. Consequently, this loss difference can be rigorously interpreted as the learning value of each feature patch, representing how much its inclusion in model training is expected to improve performance on unseen data.
However, accessing the test set for training on $D_{t}\cup D_{\mathrm{test}}$ is both prohibited and computationally impractical. To address this limitation, the validation set is adopted as a statistically sound and feasible proxy for the unseen data distribution. Accordingly, we approximate $\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{test}}\right]$ using the loss computed by a lightweight reference model trained on the validation set $D_{\mathrm{val}}$ [26]. Consequently, the resulting token weight derived in Eq. (14) provides a practical and theoretically grounded measure of each feature patch’s learning value.

### v4

# RadarPLM: Adapting Pre-trained Language Models for Marine Radar Target Detection by Selective Fine-tuning

@@ Abstract
Recent advances in pre-trained language models (PLMs) have demonstrated their capabilities in capturing universal knowledge, making them promising for radar signal processing applications. Nevertheless, directly fine-tuning PLMs on radar signals is both computationally expensive and prone to overfitting, particularly in low signal-to-clutter ratio (SCR) environments. In this paper, we propose a novel fine-tuning framework for PLM-based marine radar target detection. First, we design a lightweight adaptation module, enabling computationally efficient fine-tuning while preserving the pre-trained model’s general knowledge. Second, a novel preference-aware loss is developed to selectively optimize different feature patches based on their online-evaluated learning values, guiding the model to concentrate on those generalizable feature patterns during optimization. Finally, a binary classification head is retrained based on autoencoder network to further enhance detection performance. Experiments on real-world radar data show that the proposed RadarPLM framework yields at least a 6.35% improvement in detection performance over the existing networks under low SCR conditions. Especially, in small training samples cases, the proposed RadarPLM also achieves significant advantage over existing networks owing to the incorporation of the PLM.
@@ I Introduction
Radar-based marine small target detection faces significant challenges due to the complex sea environment, especially under low signal-to-clutter ratio (SCR) conditions. Early research, exemplified by constant false alarm rate (CFAR) detectors, hinges on the precise modeling of statistical distributions to distinguish between targets and clutter. However, accurately characterizing the non-stationary and non-Gaussian distribution of sea clutter is inherently difficult, while distribution mismatches often result in excessive false alarm rates. Furthermore, as these traditional techniques typically exploit only a single amplitude feature, their detection performance is severely compromised in low SCR scenarios.
With the development of modern wide-band radars featuring high range and velocity resolution, approaches leveraging multi-domain features of radar echoes have gained significant attention. A series of hand-crafted features [1, 2, 3, 4, 5], including the phase, Doppler, and time-frequency domain characteristics, have been devised for target detection. However, these methods rely heavily on hand-crafted heuristics and often fail to capture the nonlinear structures inherent in radar signal features.
Deep learning methods have been increasingly applied to small marine radar target detection, as they can directly and effectively learn discriminative nonlinear structures from radar signal features, offering a more flexible and powerful alternative to traditional hand-crafted approaches. Some studies use convolutional neural networks (CNNs) [6, 7, 8, 9], recurrent neural networks (RNNs) [10], and graph neural networks (GNNs) [11]. Among all the methods mentioned above, detectors based on sequence features and RNNs [10] or CNNs [9] offer an excellent trade-off between inference speed and detection performance.
Recently, the research community has experienced a paradigm shift from lightweight, task-specific neural networks toward large-scale foundation models. In particular, pre-trained large language models (LLMs), such as ChatGPT and DeepSeek, have demonstrated strong generalization capabilities across a wide range of complex tasks, including long-form text generation, mathematical reasoning, and code synthesis [12]. Motivated by these advances, an increasing body of research has extended LLMs beyond the textual domain to handle non-textual data modalities, giving rise to multimodal large language models (MLLMs). In parallel, several recent studies [13, 14, 15] have explored the adaptation of pre-trained large language models for time-series analysis and wireless signal processing. These works consistently report notable performance improvements over conventional lightweight task-specific models, particularly in scenarios characterized by limited labeled data. Building upon these developments, this work investigates the feasibility and effectiveness of leveraging PLMs as a powerful backbone network for marine radar signal processing, aiming to enhance model robustness and generalization under challenging data conditions.
@@
However, some challenges exist in adapting PLM for marine radar signal processing applications. First, owing to the large parameter scale of PLMs, their adaptation to marine radar signal processing necessitates lightweight and efficient fine-tuning strategies, particularly for deployment in resource-constrained environments. Such strategies are expected to significantly reduce the number of trainable parameters by several orders of magnitude, while preserving the pre-trained knowledge of PLMs. Second, adapting PLMs to process radar signal features in low SCR conditions often leads to severe overfitting issues. This phenomenon is fundamentally attributed to the intrinsic characteristics of radar signal features:
@@
To overcome the above-mentioned issues, including the demand for lightweight fine-tuning and the tendency of overfitting under low SCR conditions, we develop a novel framework named RadarPLM (based on GPT2, an advanced open-source PLM.). The proposed framework effectively bridges the gap between PLMs and radar target detection tasks with a lightweight adaptation module and a selective training strategy. First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens. Subsequently, a radar target detector is built upon GPT2, a highly advanced open-source PLM, and is fine-tuned through a lightweight adaptation module to efficiently align with the specific radar detection task. For model optimization, we design a preference-aware loss function, which is inspired by recent advances [16, 17]. This loss introduces a selective training strategy dynamically adjusts the learning weights of different feature patches according to their online-evaluated learning values, significantly mitigating model overfitting. Finally, the binary classification head is retrained to further improve detection performance.
Our key contributions are summarized as follows:
@@
    We propose a novel framework RadarPLM based on GPT2 model for marine radar signal processing applications. To the best of our knowledge, this work is the first to demonstrate that the recent developed PLMs can serve as a strong starting point for training and optimizing marine radar signal features.
@@
    For fine-tuning, we develop a lightweight adaptation module and design a novel selective training strategy, to reduce finetuning computational overhead and significantly mitigate the risk of model overfitting.
@@
    Extensive experiments validate that RadarPLM achieves new state-of-the-art detection performance on popular marine radar datasets, especially under practical limitations including limited training data and low SCR conditions.
Our preliminary study [18] demonstrated that fine-tuning a pre-trained GPT2 model effectively extracts discriminative information from sequence features, achieving state-of-the-art detection performance. In this paper, we further introduce a lightweight adaptation module and selective training strategy for more lightweight and generalizable adaptation.
@@ II Related Work
II-A Marine Radar Target Detection Based on Deep Learning
@@
II-B Signal Processing Models Based on Pre-trained Language Model
Recently, PLMs such as GPT2 and LLaMA2 have demonstrated exceptional capabilities in signal processing and communication tasks. For example, Zhou et al. [13] propose a unified time series analysis framework by fine-tuning frozen GPT2 [23] and achieving SOTA performance on various datasets. Liu et al. [24] propose a GPT2-empowered channel prediction framework to improve prediction accuracy. Sheng et al. [25] utilize the GPT2 model to develop an effective and robust beam prediction method. Zheng and Dai [26] propose a multi-task PLM framework to satisfy the requirements of different wireless communication tasks. IOT-LLM and Penetrative AI [27, 28] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks.
However, few of these models are specifically devised for radar signal processing applications. Moreover, recent studies [29] have raised significant concerns regarding the universal applicability and direct transferability of PLMs to various signal processing tasks. PLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks. In marine radar target detection under low SCR conditions, directly fine-tuning a PLM often leads to severe overfitting issues, limiting its parameter transfer capability. To mitigate this limitation, we introduce a preference-aware loss that selectively trains on informative feature patches.
@@ III Proposed Method
(a) IP feature
(b) DSE feature
(c) SMS feature
(d) Amp feature
(e) DP feature (Target)
(f) DP feature (Clutter)
@@ III-D1 Methodology
  • High Importance (Unlearned): If $\mathcal{L}_{\theta_{{t}}}$ is much larger than $\alpha\mathcal{L}_{\theta_{{r}}}$, the feature token is identified as containing valuable, yet unlearned patterns.
@@ III-D2 Theoretical Insight of the Effectiveness of Feature Token Importance
A. Problem Definition: Inspired by [17], we begin by formalizing an optimization problem of selecting useful feature patches for model fine-tuning. At each training step $t$, given a candidate set $B_{t}$, the objective is to select a feature patch $(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}$ whose inclusion in the training set $D_{t}$ maximally reduces the generalization loss on the unseen holdout set $D_{\mathrm{ho}}$. This can be expressed as the following optimization problem:
$\arg\min_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}-\log p\left(y^{\text{ho }}\mid F^{\text{ho }};D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right).$ (16)
@@
$-\log p\!\bigl(y^{\mathrm{ho}}\mid F^{\mathrm{ho}};D_{t}\!\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\bigr)$
$\quad=-\log\frac{p\!\bigl(y,y^{\mathrm{ho}}\mid F,F^{\mathrm{ho}},D_{t}\bigr)}{p\!\bigl(y\mid F;D_{t}\bigr)}.$ (17)
@@
$p\!\bigl(y,y^{\mathrm{ho}}\mid F,F^{\mathrm{ho}},D_{t}\bigr)\\ =p\!\bigl(y\mid F;F^{\mathrm{ho}},y^{\mathrm{ho}},D_{t}\bigr)\,p\!\bigl(y^{\mathrm{ho}}\mid F^{\mathrm{ho}};F,D_{t}\bigr)\\ =p\!\bigl(y\mid F;y^{\mathrm{ho}},F^{\mathrm{ho}},D_{t}\bigr)\,p\!\bigl(y^{\mathrm{ho}}\mid F^{\mathrm{ho}};D_{t}\bigr).$ (18)
@@
$\log p\!\left(y^{\text{ho}}\mid F^{\text{ho}};\,D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right)$ (19)
$=\log\frac{p\left(y\mid F;y^{\text{ho }},F^{\text{ho }},D_{t}\right)p\left(y^{\text{ho }}\mid F^{\text{ho }},D_{t}\right)}{p\left(y\mid F;D_{t}\right)}$
$\;\propto\;\;\log p\left(y\mid F;y^{\mathrm{ho}},F^{\mathrm{ho}},D_{t}\right)-\log p\left(y\mid F;D_{t}\right)$
$\;\propto\;\;\mathcal{L}\!\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\!\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t},D_{\text{ho}}\right].$
In summary, for a model trained on $D_{t}$, the process of identifying the feature patch that minimizes the loss on the holdout samples in Eq.(16) can be approximated by the following objective:
$\arg\max_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{ho}}\right].$ (20)
Here, $\mathcal{L}[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}]$ denotes the training loss on the patch using the current model trained on $D_{t}$, whereas $\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{ho}}\right]$ represents the excepted irreducible holdout loss obtained if the model were trained on both the training set and the unseen holdout set. This objective reveals that the difference between the training loss and the irreducible holdout loss effectively quantifies the reducible component of the generalization error contributed by each feature patch. Consequently, this loss discrepancy can be rigorously interpreted as the learning value of the patch, representing the marginal improvement in performance on unseen data expected from its inclusion in the training process.
C. Practical Implementation and Approximation: However, realizing the irreducible loss in practice for deep neural networks presents several significant challenges. We introduce the following strategic approximations to enable an efficient and tractable implementation for calculating the irreducible loss inspired by [17]:
  1. 1.
    The theoretical derivation relies on Bayesian inference and conditioning. As exact Bayesian inference is intractable in neural network, we fit the model using Stochastic Gradient Descent (SGD) instead.
  2. 2.
    In addition, we adopt the validation set ($D_{\mathrm{val}}$) as the holdout set to serve as a statistically sound proxy for the unseen data distribution.
  3. 3.
    Furthermore, since training the PLM-based model on the combined set $D_{t}\cup D_{\mathrm{val}}$ is computationally expensive, we approximate it using the loss computed by a lightweight transformer-based model (reference model) trained on $D_{\mathrm{val}}$.
D. Conclusion: Consequently, the calculated gap between the target loss $\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k})$ and the approximated irreducible loss $\mathcal{L}_{\theta_{{r}},b,k}(F^{\mathrm{fi}}_{b,k})$ (derived from the lightweight reference model) serves as the practical measure of each feature patch’s learning value, which forms the basis for the token weighting strategy defined in Eq.(14).